# xBound: Join Size Lower Bounds

### v1

@@ 5.1. Experimental Setup
We conduct the experiments on the open systems on a single node equipped with an Intel® Xeon® Gold 5318Y CPU (24 cores, 48 hyper-threads) processor and 128 GB DDR4 main memory, running Ubuntu 24.04. The experiments on Fabric Warehouse (DW) are conducted on Microsoft Fabric with a F256 capacity and version 12.0.2000.8.
Unless otherwise specified, xBound uses the following default configuration: 256 partitions, 5000 MCVs for equality predicates, and 128 histogram buckets for range predicates (the latter two are exactly those from LpBound). Moreover, we use LpBound with $\ell_{p}$ norms $\in\{\ell_{0},\ell_{1},\ell_{2},\ell_{\infty}\}$ to enable a fair comparison by applying the same set of norms to both the lower and upper bounds. [^4] In a strict sense, $\ell_{-\infty}$ is not a norm; see §2. We enable the prefix trick in LpBound, yet disable its PK-FK optimization (Ref. (Zhang et al., 2025, Sec. 5)), i.e., computing norms on two-way joins, since this incurs a high overhead, especially when building LpBound on MSSalesJoin. Both estimators store their norms into the norms tables, while xBound’s partitioned $\ell_{0}$-statistics are stored in a separate table.
Additionally, on MSSalesJoin, the partitioned $\ell_{0}$-statistics are computed in an approximate manner via a HyperLogLog sketch (Flajolet et al., 2007), as exact computation would have been too costly; we discuss the implication of this aspect in §6.
5.2. Mind the Bound
As a warm-up, we want to understand the quality of today’s system estimators, specifically, whether they lie within the safe range defined by the corresponding lower and upper bounds. The experiment is summarized in Tab. 1, where we show the percentage of bound violations per (benchmark, system) pair, when using both xBound’s lower bounds and LpBound’s upper bounds, respectively. Note that we also report the percentage of underestimates. As we see, there is a tendency to mostly underestimate, as already underlined by Leis et al. (Leis et al., 2015). It is precisely this problematic issue that xBound aims to address.
Lower bound violations.
Due to its early stage of development, DuckDB’s estimator is still imprecise. On JOBlight, 14.9% of the estimates lie below the corresponding xBound’s lower bounds. In turn, for PostgreSQL the number reads 6.9%. On the rest of the benchmarks, the numbers look more promising (from the systems’ perspective): On STATS-CEB, xBound finds 1.1% and 0.3% violated lower bounds on DuckDB and PostgreSQL, respectively. On the other side, Fabric DW stands out: Only on JOBjoin could we find a mere of 0.6% of violated lower bounds.
On the other hand, on MSSalesJoin, Fabric DW tends to overestimate (only 10.8% actual underestimates). Out of these 10.8%, only one third violates xBound’s lower bound.
Upper bound violations.
In the same Tab. 1, we also show how the systems perform relative to the corresponding upper bounds. This comparison gives us a first impression whether xBound’s lower bounds are competitive with the state-of-the-art pessimistic cardinality estimator, LpBound. On all research benchmarks, LpBound found upper bound violations across all systems, with PostgreSQL being more conservative. On MSSalesJoin, all Fabric DW’s overestimates are above the corresponding upper bounds.
5.3. Correcting Cardinality Underestimation
Discovering bound violations in estimators is not the sole goal of lower bounds (at least not that of xBound): Next, inspired by the seminal analysis by Leis et al. (Leis et al., 2015), we would like to understand whether having lower bounds can correct the infamous cardinality underestimates. Indeed, they showed more than ten years ago that systems tend to underestimate. Surprisingly, this issue is still present to this very day. Namely, we replicate their evaluation style in Tab. 2, and show the median Q-error for all (supported) subexpressions of a given benchmark, varying the number of joins. Note that, to avoid unnecessary visual clutter, we skip the number of joins from which the estimation errors are the same.
xBound-ed systems
The main observation is that xBound-ed systems, compared to the vanilla ones, have better one-sided Q-errors, i.e., xBound’s lower bounds help them approach the actual cardinality (from below). For instance, when estimating JOBlight’s subexpressions on DuckDB, the median Q-error for one join goes down from 2.34 to 2.05. On PostgreSQL, on the same number of joins, the error drops from 2.00 to 1.76. On STATS-CEB, xBound could improve DuckDB’s median Q-error from 6.72 to 6.19. Notably, on JOBjoin, xBound still manages to lower the error for three joins, namely from 2.25 to 2.06. On the rest of the benchmarks, given the limited number of lower bound violations found, we could not see any Q-error improvement.
@@ 5.4. Statistics Overhead
Relying on $\ell_{p}$-norms and min/max-values makes xBound a lightweight estimator, similar to LpBound. Indeed, in Tab. 4, we outline the overhead to build and store xBound statistics on a given DuckDB database. In terms of the database itself, we report the time needed to load it from the benchmark’s CSV files, and the resulting database file. We also report LpBound’s norms table size, and the necessary time to build the statistics.
Table 4. Space and build overhead of xBound (256 partitions).
Benchmark Database xBound LpBound
norms part. $\ell_{0}$
Size Load Size Build Size Build Size Build
[MB] [s] [MB] [s] [MB] [s] [MB] [s]
JOBlight 2646.4 50.5 5.8 26.8 21.2 27.9 2.6 6.0
JOBjoin 2646.4 50.5 0.8 24.9 ¡0.1 5.4 ¡0.1 6.2
STATS-CEB 22.9 1.2 9.4 14.3 34.3 247.5 7.6 7.8
We observe that the sizes of the norms tables of both estimators are on par. Indeed, xBound has the overhead of the prefix trick, since, while LpBound does it only for the non-filtered base tables, we have to do it in xBound even for MCV and histograms, so as to provide non-zero lower bounds. However, this extra layer of prefixes can be observed in the build time. It is worth mentioning that since xBound and LpBound share the same MCV and histogram tables, the overhead of computing them can be amortized when they are built together. In particular, because both estimators operate on the very same degree sequence (albeit sorted in different orders), one could merge the two statistics calculators to accelerate the build procedure.
Unlike LpBound, xBound also has to build, in case partitioning (§4.2) is enabled, the partitioned $\ell_{0}$-statistics. As we strive for exact distinct key counts, this will also incur an overhead, yet still comparable to that of building the actual norms. Note that the STATS-CEB benchmark has a larger number of range predicates, hence the increased overhead.
Figure 5. Average estimation time of systems and estimators for varying join sizes on all subexpressions of JOBlight.
5.5. Estimation Time
By relying on a handful of statistics, xBound’s estimation is comparable to that of traditional estimators, such as those used in DuckDB and PostgreSQL. To understand the dependency of the estimation time with respect to the number of joins, consider Fig. 5, where we plot the optimization time of the two systems, side by side with that of LpBound ($p$ = 2) and xBound. For both systems, we use the time spent to produce the plan (via EXPLAIN PLAN). For LpBound, we report only the solving time of the linear program, whereas for xBound, we measure the time taken by its two components: (i) estimating the distinct key count of the join, and (ii) applying the reverse dot-product inequalities. Note that since the lower bounds in step (i) tend to be 0 starting with 4 joins, the estimation times remain constant.
5.6. Downstream Benefits of xBound
Having shown that xBound’s lower bounds can exceed Fabric DW’s cardinality estimates, we now quantify the downstream benefit to reduce spilling—writing internal operator state to disk due to underestimating their memory consumption as consequence of cardinality underestimation. We focus on the MSSalesJoin production workload due to its realistic table/query sizes. We use Fabric DW’s query optimizer test setup to estimate operator memory consumption under the original estimates and under xBound-capped estimates (in this case, xBound uses 1024 partitions; see the ablation study in §5.7), as a function of the projected columns.
In Fig. 6, we report the resulting spilling reduction as a function of how strongly the consuming operators’ state sizes depend on the intermediate’s projected columns: from 0.3 GB when essentially none of the projected input columns contribute to state size, up to 94 GB when all columns do. The trend is linear, since the memory footprint scales with average row size times cardinality, and increasing the fraction of projected columns increases row size proportionally. Overall, even modest caps on intermediate cardinalities can translate into substantial reductions in spilling, with the realized benefit governed by the intermediate’s row width.
Figure 6. End-to-end evaluation in Fabric DW: Expected spilling reduction for MSSalesJoin when capping subexpression estimates with xBound’s lower bounds. Results vary with operator state size, which depends on how many input columns are projected.
5.7. Ablation Study
We would like to understand the effect partitioning (§4.2) has on the quality of the bounds, along with the additional optimization of norm stitching (§4.1). In the following, we took MSSalesJoin as our workload, so as to test the effect at large scale. To this end, we varied the number of partitions $\in\{1,4,16,256,1024\}$ and disabled norm stitching.
We show the results of the experiment in Fig. 7, where we plot the Q-error of xBound. [^5] We use boxen plots (Hofmann et al., 2017) since they better expose the full error distribution. The first observation is that number of partitions greatly impacts the quality of the bounds. This is due to fact that, by range partitioning, the joining keys can be better localized, hence leading to better lower bounds in xBound’s first step (§3.1). In addition, norm stitching, while orthogonal to the partitioning trick, can further boost the bounds, especially when the number of partitions is small.
Figure 7. Analyzing the impact of partitioning (§4.2) and norm stitching (§4.1) on the quality of the lower bounds on MSSalesJoin.
6. Discussion
We have demonstrated that cardinality lower bounds hold the potential to solve the long-standing cardinality underestimation problem, still present in current systems despite extensive research. In the following, we will discuss several avenues for generalizing and refining these lower bounds.
Query Type. Currently, we support only Berge-acyclic queries, with the additional requirement that the joins happen on the same join key. While this condition inherently holds for (single-key) two-way joins, e.g., the very first join in a the query plan, and, surprisingly, for the well-studied JOBlight benchmark (Kipf et al., 2019b), real-world queries often involve joins on multiple join keys.
Moreover, while rather common in established cardinality estimation benchmarks, like JOB (Leis et al., 2015) and STATS-CEB (Han et al., 2021), Berge-acyclicity no longer holds in enterprise settings to such a large extent. For instance, many two-way joins in MSSales are, in fact, not Berge-acyclic (yet, still acyclic). It is interesting to understand whether this framework can also be extended to handle cyclic queries, such as those in the Subgraph-Matching benchmark (Sun and Luo, 2020).
@@ 6. Discussion
Another intriguing extension is that to nested subqueries. While this remains an open question for LpBound, the reverse inequalities used by xBound can be extended to return lower bounds on more norms beyond $\ell_{1}$, in particular $\ell_{2}$, e.g., by using the $\ell_{4}$-norms of the input vectors. In turn, these bounds can be propagated further to subsequent subqueries.
@@
Beyond Inner Joins. To be applicable to an even more diverse workload, xBound should also support non-inner joins. We believe that the framework is flexible enough to be extended in this direction. For example, we can easily support semi-join operators, $A\ltimes B$: the lower bound on the number of distinct keys of the join $A\Join B$3.1) can be used as the prefix length when fetching the $\ell_{1}$ norm of $A$. Moreover, for outer joins, one can always return the lower bound of the (implicit) inner join. However, achieving tighter lower bounds would require more information about the number of non-joining keys, which we leave as future work.
@@
Another relevant operator is group-by. While methods exist to provide approximate estimates either via sketches (Freitag and Neumann, 2019) or learning (Kipf et al., 2019a), and even to provide upper bounds on its cardinality (Zhang et al., 2025), we are not aware of any non-probabilistic lower bounds on the cardinality of this operator.
@@
Predicate Support. While we support the standard predicate types (equality, range, conjunctions, and disjunctions), providing (good) lower bounds for string predicates, such as LIKE and REGEX, remains an open challenge. A simple technique is, naturally, sampling. The number of tuples that qualified within a sample is, indeed, a (loose) lower bound (provided that we do not extrapolate this number to the full dataset).
To generate the actual norms needed in the joins, we have to lower-bound the distinct count of the joining key (similar to the way how MCVs work, §4.3.1). The question is whether one can do better than this. We note that cardinality upper bounds can be derived using tri-grams, as hinted in SafeBound and LpBound, but applying similar techniques for lower bounds remains an open problem.
@@
Theory Advancements. As an attentive reader might have already observed, we are using a rather limited number of statistics: standard $\ell_{p}$-norms, i.e., $\{\ell_{0},\ell_{1},\ell_{2},\ell_{\infty}\}$, along with $\ell_{-\infty}$, and (partitioned) min/max values. This limited set is dictated by the class of inequalities that we are currently using. From a theory point of view, we believe that it is possible to design dot-product reverse inequalities that consider a wider range of norms, as LpBound does. Indeed, this is orthogonal to join size estimation, yet any new inequality could be easily plugged in xBound (if the corresponding norms have been computed).
@@
Deteriorating Lower Bounds. Our lower bounds rely on min/max statistics, or the so-called zonemaps (Graefe, 2009), which are readily available in database catalogs and modern file formats. While simplicity is good for real systems, collecting more granular statistics about the value distribution (not of the degree distribution, already captured by $\ell_{p}$-norms), could lead to better bounds. Indeed, their use in xBound is tightly linked to the quality of the lower bounds on the number of distinct keys in the join. Ultimately, the goal is to “locate” the joining keys more precisely, thereby mitigating the problem of deteriorating lower bounds as the number of joins increases.
xBound+HLL. To test xBound on the MSSales enterprise workload, we had to resort to approximation. Namely, instead of computing the actual numbers, we approximated the distinct key count in the partitioned statistics with HyperLogLog (Flajolet et al., 2007). While the build phase becomes faster than computing the exact distinct counts, this compromises the strict theoretical guarantees. The resulting “lower bounds” hold only in an approximate sense. We would like to understand the guarantees of xBound+HLL in a future work: Can we quantify the error of these approximate lower bounds?
@@
Data Updates. Lower bounds have an intriguing property when it comes to updates. Namely, inserts (and appends) are supported by default, since a lower bound on $|A\Join B|$ also holds when $\Delta_{A}$ is appended to $A$. This contrasts with the struggle of upper bound estimators, such as LpBound, where new data renders the previous degree norms immediately stale. Naturally, the challenge is how to incrementally maintain lower bounds on changing data. One key observation is the following: $|(A+\Delta_{A})\Join B|=|A\Join B|+|\Delta_{A}\Join B|$, due to the fact that the join size is a dot product between two vectors. Consequently, we can take the previous lower bound $|A\Join B|^{-}$, and only add the one for the newly-inserted data, $|\Delta_{A}\Join B|^{-}$. In practice, the deltas should be periodically merged to prevent storage and runtime overhead.
@@ 7. Related Work
Learned Estimators. Apart from the provable estimators in the last two decades, there has been a long line of research on learned cardinality estimation (Kipf et al., 2019b; Yang et al., 2019; Hilprecht et al., 2020; Wu et al., 2023); see, for instance, the recent analysis by Heinrich et al. (Heinrich et al., 2025). However, these estimators do not come with guarantees on their hard estimates, and can lead to arbitrarily mis-estimations when faced with complex workloads (Wehrstein et al., ). An exception is FactorJoin (Wu et al., 2023), which, indeed, provides probabilistic upper bounds.
Sketch-Based Estimators. Seeing the join size as a dot-product between the two corresponding degree vectors is not something new: Sketch-based estimators (Ganguly et al., 2004, 2005; Wang et al., 2023; Heddes et al., 2024) already use this very observation. Yet, they rely on well-established sketches, which cannot guarantee lower bounds on the actual join size; they solely return an approximation, which could, itself, violate the bounds.
Robust Query Processing. The impossibility to solve the cardinality estimation problem, let our community transition to a new paradigm, where the cardinality estimator is considered to be of secondary importance. This line of research started with Yannakakis’ algorithm (Yannakakis, 1981), and has seen a revival with a series of Dagstuhl seminars (Graefe et al., 2010), and more recently, with concrete implementations in real systems (Yang et al., 2024; Birler et al., 2024; Wang et al., 2025; Zhao et al., 2025; Bekkers et al., 2025).

### v2

@@ 5. Experiments
System Configuration. The experiments on Fabric DW are conducted on Microsoft Fabric with a F64 capacity. The standalone statistics required by xBound are built on a single node equipped with an Intel® Xeon® Gold 5318Y CPU (24 cores, 48 hyper-threads) processor and 128 GB DDR4 main memory, running Ubuntu 24.04. To this end, we use DuckDB v1.4.3.
Unless otherwise specified, xBound uses the following default configuration: 64 heavy keys, 16 light partitions, 1024 MCVs for equality predicates, and 128 histogram buckets for range predicates. We activate the following optimizations: norm stitching (§4.1), heavy keys (§3.5), which includes the extension to multiple predicates (§4.4).
5.1. Reducing Underestimation
Fabric DW’s resource estimator uses the input data volume as a proxy for estimating how many CPUs are required by the query. The data volume of an operator is computed by the number of rows times the average row size. To this end, we observed that the last join operator in the plan accounted for most of the underestimation. Since the row size, in the case of select count(*) queries, is just a constant, we directly report the estimation error on the number of rows.
Fig. 6 shows the Q-errors of the last join, whose cardinality is the query result size itself, achieved by DW and its xBound-ed variant (clipping the original estimates with the corresponding lower bounds).
We observe that (i) Fabric DW underestimates 81.5% of queries, with Q-errors reaching $5.6\cdot 10^{8}$, and (ii) its xBound-ed version reduces the median Q-error on these underestimates by 3.2x, and the Q-error’s P90 by 35.8x (it is mostly the heavily underestimated queries that xBound corrects). Since xBound can only reduce underestimation, the (mild) 18.5% overestimates of Fabric DW in this benchmark still remain.
These gains are not specific to Fabric DW. In the introductory Fig. 1, xBound reduces the Q-error of underestimates by 8.38x in DuckDB and by 2.30x in PostgreSQL on StackOverflow-CEB. On the JOBlight benchmark, only DuckDB’s median Q-error could be improved, namely by 2.45x.
5.2. Query Speedups
Correcting cardinality underestimation directly increases estimated resources for query execution in Fabric DW—thus potentially improves query execution time. Fabric DW estimates the amount of work per query stage based on the cardinalities of the contained operators (among other). Correcting underestimates via xBound translates to more CPU resources with a proportional node count rounded up to full integer—for a given query stage.
Fig. 7 shows queries with execution speedups over default Fabric DW when using resource estimates based on (i) xBound-ed cardinalities and (ii) true cardinalities. We warm up each query and then measure the time of a second query execution, limiting executions to 2h timeouts. Queries are ordered consistently with Fig. 6 and can be matched by query number. A significant number of queries show no speedup from using actual cardinalities, since doing so increases total CPU resources only minimally. This is expected based on our initial observation of long-tail CPU underestimation. However, xBound achieves significant speedup for vastly underestimated queries: 20.1x for Q90 and 3.2x for Q126. These are queries that are significantly underestimated and it shows promise that xBound may also fix the extreme CPU underestimation observed in production.
5.3. Ablation Study
We study the effects of the several optimizations introduced throughout the past sections. We study the number of partitions and the precision of the Theta sketches on the two components of xBound: the $\ell_{0}$ and the statistics of the heavy partition. Notably, each increase in the two values leads to an increased space overhead. The numbers for these have been already analyzed in §5.4.
We show the effect of these optimizations on the Q-error achieved by xBound’s lower bounds on the SO-CEB benchmark in Fig. 8. The two main observation are that the heavy partition contributes significantly to the quality of the lower bounds, and that increasing the precision of Theta sketch, especially for the $\ell_{0}$ statistics, has a better effect than increasing the number of partitions. Having a high-precision Theta sketch for heavy partition is important as well: changing the precision from $8$ to $12$ in this case improves the median Q-error from $28$ to $21$.
@@ 5.4. Statistics Overhead
Figure 7. End-to-end query speedups on Fabric DW on the SO-CEB benchmark queries. We exclude queries with no observed speedup.
Depending on the configuration, the statistics used in xBound can be considered lightweight. We report in Tab. 1 the Parquet sizes [^6] Using the DuckDB Parquet writer, when enabling zstd compression. and the time to compute all main statistics tables used in xBound: (i) the (partitioned) $\ell_{0}$-statistics stored in l0_stats, (ii) the $\ell_{p}$-norms stored in lp_stats, and (iii) the statistics related to the heavy partition, in hh_stats.
Space & Build. The overhead of the partitioned $\ell_{0}$ statistics highly depends on the number of partitions, the configuration of the Theta sketch, and of the number of join and predicate columns; the SO-CEB benchmark features 13 join columns (and equality- and range-predicates on 11 and 16 columns, respectively).
In our case, choosing 16 partitions and a precision of 8 for l0_stats results in a total size of 67.0 MB. These statistics can be built in 6.2 min. Reducing the precision to 5, which is the minimum allowed in the DataSketches library, reduces the overhead to only 13.3 MB. In contrast, the $\ell_{p}$ norms, as they do not depend on the number of partitions, take 0.2 MB. The statistics related to the hh_stats can be built fast, yet, since they require high-precision sketches (see the ablation in study in §5.3), can be rather heavyweight. A precision of 12 for this set of statistics (including the support for multiple predicates, as in §4.4) leads to 129.6 MB.
Estimation Time. Given a query on $n$ tables, we perform one Theta sketch intersection over $n$ sketches to obtain the number of distinct join keys (including a lower-bound estimate); similarly, for each of the 64 heavy keys. Our unoptimized Python prototype produces one estimate for a query in under 70 ms, with overhead largely due to pandas’ operations on data frames (including DuckDB SQL queries to fetch the sketch-based lower bounds). We expect optimized implementations to reach ¡1ms estimation time.
Figure 8. Quantifying the effect of xBound’s components (partitioned $\ell_{0}$ and heavy keys statistics) and their parameters (#parts: number of light partitions; $\theta$: ThetaSketch precision) on the Q-error of the result-size lower bounds for the supported SO-CEB queries.
6. Discussion and Open Problems
Having demonstrated the potential of cardinality lower bounds against the persistent underestimation problem, we now outline avenues for generalization, hoping to inspire both practitioners and theorists to further pursue this direction.
Query Types. Currently, xBound supports only acyclic queries whose joins share a single join key, a condition that inherently holds for two-way joins (e.g., the first join in a query plan) but is restrictive for multi-key joins common in practice. Moreover, while this type of acyclicity is prevalent in established benchmarks such as JOB (Leis et al., 2015) and StackOverflow-CEB (Han et al., 2021), it is far less so in enterprise workloads. Extending the framework to cyclic queries, such as those in the Subgraph-Matching benchmark (Sun and Luo, 2020), is an interesting open direction.
@@ 6. Discussion and Open Problems
Another intriguing extension is to nested subqueries. While this remains open for LpBound, the reverse inequalities underlying xBound can be generalized to yield lower bounds on norms beyond $\ell_{1}$, in particular $\ell_{2}$, e.g., via the $\ell_{4}$-norms of the input vectors, enabling propagation to subsequent subexpressions.
@@
Beyond Inner Joins. xBound is able to support non-inner joins to handle a more diverse workload. The framework is flexible enough for this: for example, semi-joins $A\ltimes B$ can be supported by using the lower bound on the number of distinct join keys (§3.1, §3.2) as the prefix length when fetching the $\ell_{1}$ norm of $A$. For outer joins, one can always return the inner-join lower bound; tighter bounds would require information about non-joining keys, which we leave as future work.
Table 1. Space and build overhead (16 partitions). We vary the ThetaSketch precision ($\theta$) used by the statistics related to $\ell_{0}$ (l0_stats) and the heavy partition (hh_stats).
Table Purpose $\theta$ Size [MB] Build [s]
l0_stats Partitioned $\ell_{0}$-stats 5 13.3 360.1
8 67.0 373.9
lp_stats $\ell_{p}$-norms 0.2 163.0
hh_stats Heavy partition stats 8 22.6 18.1
12 129.6 26.5
@@
Another relevant operator is group-by, for which approximate estimates exist via sketches (Freitag and Neumann, 2019) or learning (Kipf et al., 2019a), as well as upper bounds (Zhang et al., 2025).
@@
Predicate Support. While we support standard predicate types (equality, range, conjunctions, and disjunctions), providing good lower bounds for string predicates such as LIKE and REGEX remains open. A simple approach is sampling: the number of qualifying tuples in a sample is a (loose) lower bound, provided we do not extrapolate to the full dataset. To generate the norms needed for joins, we must lower-bound the distinct count of the join key (as with MCVs, §4.3.1); whether one can do better is an open question. We note that cardinality upper bounds can be derived via tri-grams, as hinted in SafeBound and LpBound, but extending such techniques to lower bounds remains unexplored.
@@
Theory Advancements. As the reader may have observed, xBound relies on a limited set of statistics: $\{\ell_{0},\ell_{1},\ell_{2},\ell_{\infty}\}$, $\ell_{-\infty}$, and (partitioned) min/max values, dictated by the class of inequalities we currently use. We believe it is possible to design reverse inequalities over a wider range of norms, as LpBound does; any such inequality could be readily plugged into xBound once the corresponding norms are available.
@@
Deteriorating Lower Bounds. Our exact lower bounds rely on min/max statistics (zonemaps (Graefe, 2009)), readily available in database catalogs and modern file formats. While this simplicity suits real systems, more granular statistics about the value distribution (beyond the degree distribution already captured by $\ell_{p}$-norms) could yield tighter bounds. Their role in xBound is tied to the quality of lower bounds on the number of distinct join keys; ultimately, the goal is to locate joining keys more precisely, mitigating the deterioration of lower bounds as the number of joins increases.
Probabilistic Lower Bounds. To be able to run xBound on large datasets, we had to resort to probabilistic lower bounds. This affected the $\ell_{0}$ statistics (computed at 99% confidence) and multi-predicate support for heavy partitions. Notably, the inner-product reverse inequalities remained exact. An open question is understanding the joint guarantee when intersecting predicate-level Theta sketches for all heavy keys: while each key’s degree is taken at 99% confidence, the overall guarantee degrades as these operations are performed jointly across all heavy keys.
@@
Data Updates. Lower bounds have an intriguing property regarding updates: inserts are supported by default, since a lower bound on $|A\Join B|$ remains valid when $\Delta_{A}$ is appended to $A$, unlike upper bound estimators such as LpBound, where new data immediately renders degree norms stale. For incremental maintenance, we observe that $|(A+\Delta_{A})\Join B|=|A\Join B|+|\Delta_{A}\Join B|$, as the join size is an inner product. Thus, we can reuse the previous lower bound $|A\Join B|^{-}$ and add only $|\Delta_{A}\Join B|^{-}$ for the new data, periodically merging deltas to limit storage and runtime overhead.
@@ 7. Related Work
Learned Estimators. Apart from the provable estimators in the last two decades, there has been a long line of research on learned cardinality estimation (Kipf et al., 2019b; Yang et al., 2019; Hilprecht et al., 2020; Wu et al., 2023); see, for instance, the recent analysis by Heinrich et al. (Heinrich et al., 2025). However, these estimators do not come with guarantees on their hard estimates, and can lead to arbitrarily mis-estimations when faced with complex workloads (Wehrstein et al., [n. d.]). An exception is FactorJoin (Wu et al., 2023), which, indeed, provides probabilistic upper bounds.
Sketch-Based Estimators. Seeing the join size as an inner product between the two corresponding degree vectors is not something new: Sketch-based estimators (Ganguly et al., 2004, 2005; Wang et al., 2023; Heddes et al., 2024) already use this very observation. Yet, they rely on well-established sketches, which cannot guarantee lower bounds on the actual join size; they solely return an approximation, which could, itself, violate the bounds.
Robust Query Processing. The impossibility to solve the cardinality estimation problem let our community transition to a new paradigm, where the cardinality estimator is considered to be of secondary importance. This line of research started with Yannakakis’ algorithm (Yannakakis, 1981), and has seen a revival with a series of Dagstuhl seminars (Graefe et al., 2010), and more recently, with concrete implementations in modern database systems (Yang et al., 2024; Birler et al., 2024; Wang et al., 2025; Zhao et al., 2025; Bekkers et al., 2025).